OpenStudy (anonymous):

Simplify the expression (k^1/2)^-9 _____________ (k^4)^1/8

OpenStudy (anonymous):

\[\frac{ k ^{-9/2} }{ k ^{4/8} } = \frac{ k ^{-9/2} }{ k ^{1/2} }=\frac{ (k ^{1/2})^{-9} }{ k ^{1/2} }\]

OpenStudy (anonymous):

Thanks Dharmaputra.